Circular No. 5655

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

GALACTIC CENTER I. F. Mirabel, Service d’Astrophysique, Centre d’Etudes de Saclay; L.F.

Rodriguez, Universidad Nacional Autonoma de Mexico; and B. Cordier, J. Paul, and F. Lebrun, Service d’Astrophysique, Centre d’Etudes de Saclay, communicate: “In coordination with SIGMA, we are carring out Very Large Array (VLA) radio observations of 1E 1740.7–2942 and GRS 1758–258, the two strongest persistent soft γ-ray sources in the galactic center region. Within the SIGMA error box for GRS 1758–258, there is a radio source that has shown a flux variation in the same sense as the 40- to 150-keV photon counts of SIGMA. In 1992 January-April, the 6-cm flux from this source was below our detection limit of 0.07 mJy. On Sept. 10 it appeared with a flux of 0.48 ± 0.05 mJy, and on Sept. 26 it had a flux of 0.58 ± 0.06 mJy. A similar trend was observed by SIGMA in soft γ-rays, since GRS 1758–258 was not detected during 1992 February-April above the 1σ detection limit of 8 mCrab, whereas in 1992 September it reappeared with a 40- to 150-keV flux of 35 ± 12 mCrab (Cordier et al., Compton Symp. 1992). This time-variable radio source is likely to be the radio counterpart of GRS 1758–258. Its position on Sept. 10 was α = 17^{h}58^{m}06^{s}.3, δ = −25^{o}44′34′′.0 ± 2′′ (equinox 1950.0), which is at the center of the double-sided radio jets reported by Rodriguez, Mirabel, and Marti (1992, Ap.J. 401, L15).”

I. F. Mirabel and P. A. Duc, Service d’Astrophysique, Centre d’Etudes de Saclay, report: “We are searching for infrared counterparts of 1E 1740.7– 2942 and GRS 1758–258 with the 2.2-m telescope of the European Southern Observatory. H, J, and K images taken on June 16 and 17 reveal no infrared counterparts down to mag K = 17 within 2′′ of the VLA compact variable radio sources associated with the soft γ-ray sources. For 1E 1740.7– 2942 our result is comparable to that obtained by Prince and Skinner (IAUC 5252), and more recently by Djorgovski et al. (IAUC 5596) on 1992 July 21. If the optical absorption to these sources is < 50 mag, the magnitude limits at 2.2 μm imply that the binary companions of these black-hole candidates are not massive stars with optical luminosities brighter than M_{v} = –3 (as in Cygnus X-1).”

NOVA PUPPIS 1991 Photometry (± 0.02) by A. C. Gilmore, Mount John University Ob-

servatory (cf. IAUC 5628): Oct. 15.64 UT, V = 11.94, B–V = +0.02, U–B = –1.11, V–R = +0.38.

1992 November 18 Daniel W. E. Green